A jeep has to reach one peak at 89.6 high to another 70 m high .the horizontl distance of seperation is 20 m .the minimum velocity  with which it has to jump

Grade:12

2 Answers

JOSEPH KIMUTAI
5 years ago
height h=89.6-70=19.6m
range,R=20m
but R=ut
then h=0.5gt^2
t^2=h/0.5g 
t^2=19.6/0.5*10
t^2=3.92
t=1.9799s
R=ut
u=R/t
=20/1.9799
10.1m/s
abc
5 years ago
 
height h=89.6-70=19.6m
range,R=20m
but R=ut
then h=0.5gt^2
t^2=h/0.5g 
t^2=19.6/0.5*10
t^2=3.92
t=1.9799s
R=ut
u=R/t
=20/1.9799
10.1m/s